a company ships computer components in boxes that contain 20 items. Assume that the probablity of a defective computer components is 0.2.Use the geometric variance to find the variance of defective parts.
Answer by Edwin McCravy(20055)

Variance = (standard deviation)² = s² = n·p·(1-p) = 20·(0.2)·(1-0.2) = 20·(0.2)·(0.8) = 3.2